What are the center radius form of a circle defined by x²+y²–2x+4y–1=0

1
Expert's answer
2022-01-24T18:29:49-0500
"x^2+y^2-2x+4y-1=0"

"(x^2-2x+1)+(y^2+4y+4)=6"

"(x-1)^2+(y+2)^2=(\sqrt{6})^2"

The center-radius form of the circle equation is

"(x-1)^2+(y+2)^2=(\sqrt{6})^2,"

with the center at "(1, -2)" and the radius "\sqrt{6}."